# Impact of Valley Degeneracy on Thermoelectric Properties of Zigzag Graphene Nanoribbons with Staggered Sublattice Potentials and Transverse Electric Fields

### v1

@@ III.2 Transmission Coefficient of 4-AB/4-C/4-AB under Transverse Electric Fields
As observed in Fig. 3, the curvature of the conduction (valence) subband in 4-AB/4-C/4-AB ZGNRs can be adjusted by a transverse electric field. This indicates that the effective masses of electrons (holes) and the density of states near the subband edge depend on the applied voltage strength ($V_{y}$). It is expected that the ${\cal T}_{LR}(\varepsilon)$ curves are also significantly influenced by these transverse electric fields. To examine the impact of $V_{y}$, we present the calculated ${\cal T}_{LR}(\varepsilon)$ curves for 4-AB/4-C/4-AB ZGNR segments across various $V_{y}$ values in Fig. 7. The spectra of ${\cal T}_{LR}(\varepsilon)$ depend not only on the magnitude of the electric fields but also on their direction. Moreover, the shapes of the ${\cal T}_{LR}(\varepsilon)$ curves vary distinctly with each $V_{y}$. In reference [Whitney, 27], it was noted that efficient heat engines with optimized power outputs can be realized when the ${\cal T}_{LR}(\varepsilon)$ curve exhibits an ideal rectangular shape. However, such an ideal rectangular curve (${\cal T}_{LR}(\varepsilon)=N_{v}$ for $\varepsilon<\varepsilon_{HO}$, where $N_{v}$ represents valley degeneracy) exists only in infinitely long 1D systems without defects [Areshkin, 60-Matsuda, 64].
@@ III.3 Thermoelectric Properties of Finite 4-AB/4-C/4-AB ZGNR Segments
The transmission coefficients presented in Fig. 6 and Fig. 7 reveal the electrical conductance $G_{e}(\mu)$ of 4-AB/4-C/4-AB ZGNR segments at zero temperature, given by $G_{e}(\mu)=G_{0}{\cal T}_{LR}(\mu)$, where $G_{0}=\frac{2e^{2}}{h}$ is the quantum conductance. For ZGNRs with two valley degeneracy, the ideal $G_{e}(\mu)$ reaches $\frac{4e^{2}}{h}$[XieXC, 79]. To illustrate the contact effects on the thermoelectric properties of 4-AB/4-C/4-AB ZGNRs, we plot the curves of electrical conductance ($G_{e}$), Seebeck coefficient ($S$), power factor ($PF=S^{2}G_{e}$), and figure of merit ($ZT$) at $T=324$ K in Fig. 8. Throughout this article, we use specific constants: $G_{0}=2e^{2}/h=77.5~{}\mu S$ for electrical conductance, $k_{B}/e=86.25\mu V/K$ for the Seebeck coefficient, and $2k_{B}^{2}/h=0.575pW/K^{2}$ for the power factor.
In Fig.8(a), the curves of $G_{e}(\mu)$ at finite temperature are influenced by the averaged ${\cal T}_{LR}(\varepsilon)$ curves, with the averaging range of $\varepsilon$ determined by $\frac{1}{4k_{B}T\cosh^{2}((\varepsilon-\mu)/(2k_{B}T))}$. At low temperatures, states near $\mu$ dominate charge transport. Fig. 8(b) shows that the maximum Seebeck coefficient occurs at small $\Gamma_{t}=0.27$ eV, although the six $S$ curves nearly coincide in the range $\varepsilon_{HO}<\mu<-0.45$ eV. In Fig. 8(c) and 8(d), the maximum power factor $PF_{max}=1.16$ and figure of merit $ZT_{max}=3.46$ are observed at $\Gamma_{t}=0.81$ eV. The figure of merit $ZT$ is calculated using $ZT=\frac{S^{2}G_{e}T}{\kappa_{e}+\kappa_{ph}}$, where $\kappa_{ph}$ represents the phonon thermal conductance of the 4-AB/4-C/4-AB structure, approximated here as $\kappa_{ph}=F_{s}\kappa_{GNR}$ with $\kappa_{GNR}=\frac{\pi^{2}k_{B}^{2}T}{3h}$. Here, $F_{s}=0.1$ accounts for substrate effects, reducing $\kappa_{ph}$ as shown theoretically for ZGNRs with boron nitride interfaces [TangXF, 80Algharagholy, 83]. Note that $PF_{max}$ and $ZT_{max}$ for $\Gamma_{t}=0.81$ eV occur at chemical potentials $\mu=-0.823$ eV and $\mu=-0.787$ eV, respectively. $ZT$ values exceeding three are found only within a very small range of $\mu$. The optimized chemical potential for maximizing the power factor is $21$ meV above $\epsilon_{HO}=-0.844$ eV, known as thermionic-assisted transport (TAT). Experimentally, achieving such optimized positions of electrode chemical potentials depicted in Fig. 8 presents a challenge due to difficulties in precise modulation[Ranganayakulu, 36]. Introducing transverse electric fields may offer practical means to adjust the separation between $\varepsilon_{HO}$ and $\mu$ levels.
@@
Unlike electrical conductance, Seebeck coefficients exhibit a temperature-dependent bipolar behavior. The positive and negative signs reflect competition between electrons and holes, corresponding to occupied states above $\mu$ and unoccupied states below $\mu$, respectively. As temperature increases beyond $k_{B}T>10$ meV, positive $S$ values indicate dominance of hole carriers. In this temperature range, higher temperatures lead to larger Seebeck coefficients, highlighting temperature-dependent electron-hole asymmetry in DBT when $\mu$ approaches or falls below $\varepsilon_{HO}$. For instance, the curve with $\Delta=-20$ meV and $V_{y}=-0.9$V exhibits a maximum $S$ value of $138\mu V/K$ at $k_{B}T=80$meV, which is higher than the $120\mu V/K$ observed for $\Delta=-9$ meV and $V_{y}=0.9$ V at the same temperature. Consequently, the maximum power factor $PF_{max}=1.21\cdot\frac{2k_{B}^{2}}{h}$ at $k_{B}T=80$ meV occurs for $\Delta=-20$ meV and $V_{y}=-0.9$ V.
The variations in chemical potential depicted in Fig. 9 correspond to behaviors observed in codoped Germanium Telluride ($Ge_{1-x-y}Sb_{x}Bi_{y}Te$) thermoelectric materials. It is noteworthy that the electronic structures of $GeTe$ exhibit multiple quasi-1D valley structures near the Fermi energy [Ranganayakulu, 36]. The behaviors of electrical conductance, Seebeck coefficient, and power factor with respect to temperature in Fig. 9(a), 9(b), and 9(c) effectively elucidate the thermoelectric properties of these materials within the temperature range of $10$ meV $<k_{B}T<70$ meV. The arch-shaped transmission coefficient curve provides insights into the potential mechanism underlying the thermoelectric properties of $GeTe$ materials with multiple valley degeneracies.
@@ IV Conclusion
In 4-AB/4-C/4-AB ZGNR segments, the $|\Psi_{\ell,j}(E)|^{2}$ of level degeneracy are delocalized and asymmetrically distributed along the transverse direction. Their transmission coefficient (${\cal T}_{LR}(\varepsilon)$) spectra are significantly influenced by contact properties associated with electrode materials. The shape of the ${\cal T}_{LR}(\varepsilon)$ curve can be manipulated not only by the magnitude but also by the direction of the applied transverse electric field, characteristics crucial for applications in managing electron heat currents [Martinez, 84Kuo5, 86].
The thermoelectric efficiency of 4-AB/4-C/4-AB ZGNRs with optimized electrical power output favors thermionic-assisted transport (TAT). Level degeneracy enhances electrical conductance while maintaining Seebeck coefficients, thereby enhancing the power factor. In TAT, $ZT$ can also be improved due to the enhanced power factor under conditions where $\kappa_{ph}\gg\kappa_{e}$. Conversely, in direct ballistic transport (DBT), the power factor is reduced due to the small Seebeck coefficient at low temperatures. As $\mu$ approaches and falls below $\varepsilon_{HO}$, significant enhancements in the Seebeck coefficient at high temperatures and modest reductions in electrical conductance lead to enhanced power factors. This phenomenon results from increased electron-hole asymmetrical range with rising temperature in 4-AB/4-C/4-AB ZGNRs. The results in Fig. 9(a), 9(b), and 9(c) within the range of $k_{B}T=20$ meV to $k_{B}T=70$ meV provide valuable insights into the thermoelectric properties observed in materials like Germanium Telluride [Ranganayakulu, 36].
@@ Acknowledgments
This work was supported by the National Science and Technology Council, Taiwan under Contract No. MOST 107-2112-M-008-023MY2.
@@ Appendix A Electronic Band Structures of ZGNRs with w-AB/n-C/w-AB Scenario
To further understand the electronic properties depicted in Fig. 2(a), we present the electronic structures of 4-AB/4-C/4-AB for various $\delta$ values in Fig. A.1. In Fig. A.1(a) with $\delta=0$, the electronic structure of 4-AB/4-C/4-AB exhibits metallic properties of ZGNRs, characterized by flat bands near the CNP [Nakada, 66-Wakabayashi, 67]. In Fig. A.1(b), with $\delta=0.45$ eV, only the frontier conduction and valence subbands are perturbed. Other subbands remain largely unaffected, and a small gap separates the conduction subband from the valence subband due to the presence of staggered sublattice potential. Additionally, the originally flat band observed at $\delta=0$ is slightly distorted.

### v3

@@ III.2 Transmission Coefficient of 4-AB/4-C/4-AB under Transverse Electric Fields
As observed in Fig. 3, the curvature of the conduction (valence) subband in 4-AB/4-C/4-AB ZGNRs can be adjusted by a transverse electric field. This indicates that the effective masses of electrons (holes) and the density of states near the subband edge depend on the applied voltage strength ($V_{y}$). It is expected that the ${\cal T}_{LR}(\varepsilon)$ curves are also significantly influenced by these transverse electric fields. To examine the impact of $V_{y}$, we present the calculated ${\cal T}_{LR}(\varepsilon)$ curves for 4-AB/4-C/4-AB ZGNR segments across various $V_{y}$ values at $\Gamma_{t}=2.7$ eV in Fig. 7. According to the results of Fig. 6, the spectra of the transmission coefficient curve at $\Gamma_{t}=2.7$ eV can reach two, illustrating the valley degeneracy. Therefore, we set $\Gamma_{t}=2.7$ eV in Fig.7. The spectra of ${\cal T}_{LR}(\varepsilon)$ depend not only on the magnitude of the electric fields but also on their direction. Moreover, the shapes of the ${\cal T}_{LR}(\varepsilon)$ curves vary distinctly with each $V_{y}$. In reference [Whitney, 32], it was noted that efficient heat engines with optimized power outputs can be realized when the ${\cal T}_{LR}(\varepsilon)$ curve exhibits an ideal rectangular shape. However, such an ideal rectangular curve (${\cal T}_{LR}(\varepsilon)=N_{v}$ for $\varepsilon<\varepsilon_{HO}$, where $N_{v}$ represents valley degeneracy) exists only in infinitely long 1D systems without defects [Areshkin, 68-Matsuda, 72].
@@ III.3 Thermoelectric Properties of Finite 4-AB/4-C/4-AB ZGNR Segments
The transmission coefficients presented in Fig. 6 and Fig. 7 reveal the electrical conductance $G_{e}(\mu)$ of 4-AB/4-C/4-AB ZGNR segments at zero temperature, given by $G_{e}(\mu)=G_{0}{\cal T}_{LR}(\mu)$, where $G_{0}=\frac{2e^{2}}{h}$ is the quantum conductance. For ZGNRs with two valley degeneracy, the ideal $G_{e}(\mu)$ reaches $\frac{4e^{2}}{h}$[XieXC, 87]. To illustrate the contact effects on the thermoelectric properties of 4-AB/4-C/4-AB ZGNRs, we plot the curves of electrical conductance ($G_{e}$), Seebeck coefficient ($S$), power factor ($PF=S^{2}G_{e}$), and figure of merit ($ZT$) at $T=324$ K in Fig. 8. Throughout this article, we use specific constants: $G_{0}=2e^{2}/h=77.5~{}\mu S$ for electrical conductance, $k_{B}/e=86.25\mu V/K$ for the Seebeck coefficient, and $2k_{B}^{2}/h=0.575pW/K^{2}$ for the power factor, based on of Eq.(2).
In Fig.8(a), the curves of $G_{e}(\mu)$ at finite temperature are influenced by the averaged ${\cal T}_{LR}(\varepsilon)$ curves, with the averaging range of $\varepsilon$ determined by $\frac{1}{4k_{B}T\cosh^{2}((\varepsilon-\mu)/(2k_{B}T))}$. At low temperatures, states near $\mu$ dominate charge transport. Fig. 8(b) shows that the maximum Seebeck coefficient occurs at small $\Gamma_{t}=0.27$ eV, although the six $S$ curves nearly coincide in the range $\varepsilon_{HO}<\mu<-0.45$ eV. In Fig. 8(c) and 8(d), the maximum power factor $PF_{max}=1.16$ and figure of merit $ZT_{max}=3.46$ are observed at $\Gamma_{t}=0.81$ eV. The figure of merit $ZT$ is calculated using $ZT=\frac{S^{2}G_{e}T}{\kappa_{e}+\kappa_{ph}}$, where $\kappa_{ph}$ represents the phonon thermal conductance of the 4-AB/4-C/4-AB structure, approximated here as $\kappa_{ph}=F_{s}\kappa_{GNR}$ with $\kappa_{GNR}=\frac{\pi^{2}k_{B}^{2}T}{3h}$. Here, $F_{s}=0.1$ accounts for substrate effects, reducing $\kappa_{ph}$ as shown theoretically for ZGNRs with boron nitride interfaces [TangXF, 88Algharagholy, 91]. Note that $PF_{max}$ and $ZT_{max}$ for $\Gamma_{t}=0.81$ eV occur at chemical potentials $\mu=-0.823$ eV and $\mu=-0.787$ eV, respectively. $ZT$ values exceeding three are found only within a very small range of $\mu$. The optimized chemical potential for maximizing the power factor is $21$ meV above $\epsilon_{HO}=-0.844$ eV, known as thermionic-assisted transport (TAT). Experimentally, achieving such optimized positions of electrode chemical potentials depicted in Fig. 8 presents a challenge due to difficulties in precise modulation[Ranganayakulu, 41]. Introducing transverse electric fields may offer practical means to adjust the separation between $\varepsilon_{HO}$ and $\mu$ levels.
@@
Unlike electrical conductance, Seebeck coefficients exhibit a temperature-dependent bipolar behavior. The positive and negative signs reflect competition between electrons and holes, corresponding to occupied states above $\mu$ and unoccupied states below $\mu$, respectively. As temperature increases beyond $k_{B}T>10$ meV, positive $S$ values indicate dominance of hole carriers. In this temperature range, higher temperatures lead to larger Seebeck coefficients, highlighting temperature-dependent electron-hole asymmetry in DBT when $\mu$ approaches or falls below $\varepsilon_{HO}$. For instance, the curve with $\Delta=-20$ meV and $V_{y}=-0.9$V exhibits a maximum $S$ value of $138\mu V/K$ at $k_{B}T=80$meV, which is higher than the $120\mu V/K$ observed for $\Delta=-9$ meV and $V_{y}=0.9$ V at the same temperature. Consequently, the maximum power factor $PF_{max}=1.21\cdot\frac{2k_{B}^{2}}{h}$ at $k_{B}T=80$ meV occurs for $\Delta=-20$ meV and $V_{y}=-0.9$ V. To demonstrate that the results in Fig. 9 are not significantly affected by a larger $N_{z}$, we compare the transmission coefficients at $N_{z}=81$ and $N_{z}=101$(with $L_{a}=12.3$ nm) in Appendix B. We find that these two curves show very little difference. This indicates that the $N_{z}=81$ used to calculate the transmission coefficients in Figs. 6 and 7 is sufficiently converged for $\Gamma_{t}=2.7$ eV.
The variations in chemical potential depicted in Fig. 9 correspond to behaviors observed in co-doped Germanium Telluride ($Ge_{1-x-y}Sb_{x}Bi_{y}Te$) thermoelectric materials. It is noteworthy that the electronic structures of $GeTe$ exhibit multiple quasi-1D valley structures near the Fermi energy [Ranganayakulu, 41]. The behaviors of electrical conductance, Seebeck coefficient, and power factor with respect to temperature in Fig. 9(a), 9(b), and 9(c) effectively elucidate the thermoelectric properties of these materials within the temperature range of $10$ meV $<k_{B}T<70$ meV. The arch-shaped transmission coefficient curve provides insights into the potential mechanism underlying the thermoelectric properties of $GeTe$ materials with multiple valley degeneracies.
@@ IV Conclusion
In 4-AB/4-C/4-AB ZGNR segments, the $|\Psi_{\ell,j}(E)|^{2}$ of level degeneracy are delocalized and asymmetrically distributed along the transverse direction. Their transmission coefficient (${\cal T}_{LR}(\varepsilon)$) spectra are significantly influenced by contact properties associated with electrode materials. The shape of the ${\cal T}_{LR}(\varepsilon)$ curve can be manipulated not only by the magnitude but also by the direction of the applied transverse electric field, characteristics crucial for applications in managing electron heat currents [Martinez, 92Kuo5, 94].
The thermoelectric efficiency of 4-AB/4-C/4-AB ZGNRs with optimized electrical power output favors thermionic-assisted transport (TAT). Level degeneracy enhances electrical conductance while maintaining Seebeck coefficients, thereby enhancing the power factor. In TAT, $ZT$ can also be improved due to the enhanced power factor under conditions where $\kappa_{ph}\gg\kappa_{e}$. Conversely, in direct ballistic transport (DBT), the power factor is reduced due to the small Seebeck coefficient at low temperatures. As $\mu$ approaches and falls below $\varepsilon_{HO}$, significant enhancements in the Seebeck coefficient at high temperatures and modest reductions in electrical conductance lead to enhanced power factors. This phenomenon results from increased electron-hole asymmetrical range with rising temperature in 4-AB/4-C/4-AB ZGNRs. The results in Fig. 9(a), 9(b), and 9(c) within the range of $k_{B}T=20$ meV to $k_{B}T=70$ meV provide valuable insights into the thermoelectric properties observed in materials like Germanium Telluride [Ranganayakulu, 41].
@@ Acknowledgments
This work was supported by the National Science and Technology Council, Taiwan under Contract No. MOST 107-2112-M-008-023MY2. We acknowledge the use of the computing facility provided by the Research Center for Applied Sciences, Academia Sinica, Taiwan.
@@ Appendix A Electronic Band Structures of ZGNRs with w-AB/n-C/w-AB Scenario
To further understand the electronic properties depicted in Fig. 2(a), we present the electronic structures of 4-AB/4-C/4-AB for various $\delta$ values in Fig. A.1. In Fig. A.1(a) with $\delta=0$, the electronic structure of 4-AB/4-C/4-AB exhibits metallic properties of ZGNRs, characterized by flat bands near the CNP [Nakada, 74-Wakabayashi, 75]. In Fig. A.1(b), with $\delta=0.45$ eV, only the frontier conduction and valence subbands are perturbed. Other subbands remain largely unaffected, and a small gap separates the conduction subband from the valence subband due to the presence of staggered sublattice potential. Additionally, the originally flat band observed at $\delta=0$ is slightly distorted.
@@
Appendix B Transmission coefficients of 4-AB/4-C/4-AB ZGNRs under transverse electric fields
The thermoelectric quantities in Fig. 8 and Fig. 9 are calculated using $N_{z}=81$. We aim to demonstrate that the results in these figures do not depend on the length $L_{z}$. In Fig. B.1, we present the transmission coefficient for 4-AB/4-C/4-AB ZGNRs with $N_{z}=101$ (where $L_{z}=12.3$ nm). The transmission coefficient curve exhibits electron-hole symmetry, and the solid curve ($N_{z}=101$) is difficult to distinguish from the dashed curve ($N_{z}=81$). This implies that the length of the 4-AB/4-C/4-AB ZGNR segment with $N_{z}=81$ is sufficiently long at $\Gamma_{t}=2.7$ eV.
Figure B.1: Transmission coefficient of 4-AB/4-C/4-AB ZGNRs with $\delta=0.9$ eV, $V_{y}=-0.9$ V, and $\Gamma_{t}=2.7$ eV.